Charlie invested $2,000 at the end of 2015, $2,500 at the end of 2016 and $550...

  1. Charlie invested $2,000 at the end of 2015, $2,500 at the end of 2016 and $550 at the end of 2017.  She earned 8% on her investment.  How much money did she have at the end of 2017?
Answer #1

This question requires application of basic time value of money function, according to which

FV = PV * (1 + r)n

Now, for this question,

Total FV = 2000 * (1 + 8%)2 + 2500 * (1 + 8%)1 + 550 * (1 + 8%)0

Total FV = 2,332.80 + 2,700 + 550

Total FV = $5,582.80